A chef is using a mixture of two brands of Italian dressing. The first brand contains 7% vinegar, and the second brand contains 11% vinegar. The chef wants to make 320 milliliters of a dressing that is 8% vinegar. How much of each brand should she use?

Answer by checkley77(12844)

.11X+.07(320-X)=.08*320
.11X+22.4-.07X=25.6
.04X=25.6-22.4
.04X=3.2
X=3.2/.04
X=80 ML OF THE 11% VINEGAR.
320-80=240 ML OF 7% VINEGAR.
PROOF:
.11*80+.07*240=25.6
8.8+16.8=25.6
25.6=25.6